Rank these metals on the basis of their cutoff frequency.
Platinium 6.4 eV, Beryllium 5.0 eV, Magnesium 3.7 eV, Potassium 2.3 eV, Cesium 2.1 eV
Answer #1

The relationship between the work function and cutoff frequency is The cutoff frequency is directly proportional to work func
Answer #2
Work Functions for Photoelectric Effect

Element Work Function(eV)
Aluminum
4.08
Beryllium
5.0
Cadmium
4.07
Calcium
2.9
Carbon
4.81
Cesium
2.1
Cobalt
5.0
Copper
4.7
Gold
5.1
Iron
4.5
Lead
4.14
Magnesium
3.68
Mercury
4.5
Nickel
5.01
Niobium
4.3
Potassium
2.3
Platinum
6.35
Selenium
5.11
Silver
4.73
Sodium
2.28
Uranium
3.6
Zinc
4.3
Answer #3
we know that
E = h * f
where E is energy,h = 6.63 * 10^-34 J-s and f is frequency
or f = E/h